# Copula‐Stein Discrepancy: A Generator‐Based Stein Operator for Archimedean Dependence

### v1

@@ 6.2 Finite‐Sample Convergence and Optimality (Detailed Proof)
Since $\mathbb{E}[\widehat{\mathrm{CSD}}_{n}^{2}]=O(n^{-1})$, it follows that $\widehat{\mathrm{CSD}}_{n}^{2}=O_{P}(n^{-1})$. Finally, taking square roots (a continuous mapping) yields $\widehat{\mathrm{CSD}}_{n}=O_{P}(n^{-1/2})$.
6.3 Sensitivity to Tail Dependence
Theorem 5 (Tail-dependence sensitivity)
$\mathrm{CSD}_{C_{2}}(C_{1})\;>\;0.$
Proof. We write $C_{0}\equiv C_{2}$ for brevity. We prove the lower tail case. The idea is to solve the Stein equation for a lower-tail indicator and evaluate the resulting witness under $C_{1}$.
Fix a small $\varepsilon\in(0,1)$. Let $B_{\varepsilon}\equiv[0,\varepsilon]^{2}$. Let $\rho_{\delta}$ be a standard $C^{\infty}$ mollifier on $\mathbb{R}^{2}$ with support in a ball of radius $\delta$. Define $h_{\varepsilon,\delta}\equiv\mathbf{1}_{B_{\varepsilon}}*\rho_{\delta}$ restricted to $(0,1)^{2}$, and then center it under $C_{0}$ by $\bar{h}_{\varepsilon,\delta}\equiv h_{\varepsilon,\delta}-\mathbb{E}_{C_{0}}[h_{\varepsilon,\delta}]$. The function $\bar{h}_{\varepsilon,\delta}$ is bounded Lipschitz uniformly in $\delta$ with unit sup norm and Lipschitz constant of order $\delta^{-1}$.
By the Stein equation construction in the proof of Theorem 3, there exists $g_{\varepsilon,\delta}\in\mathcal{H}^{2}$ with
$\mathcal{T}_{C_{0}}g_{\varepsilon,\delta}=\bar{h}_{\varepsilon,\delta}\qquad\text{and}\qquad\|g_{\varepsilon,\delta}\|_{\mathcal{H}^{2}}\leq C\|\bar{h}_{\varepsilon,\delta}\|_{\mathrm{BL}}\leq C^{\prime}\delta^{-1}.$
Therefore,
$\mathrm{CSD}_{C_{2}}(C_{1})\ \geq\ \frac{\mathbb{E}_{C_{1}}[\mathcal{T}_{C_{0}}g_{\varepsilon,\delta}]}{\|g_{\varepsilon,\delta}\|_{\mathcal{H}^{2}}}\ =\ \frac{\mathbb{E}_{C_{1}}[\bar{h}_{\varepsilon,\delta}]}{\|g_{\varepsilon,\delta}\|_{\mathcal{H}^{2}}}\ \geq\ \frac{\mathbb{E}_{C_{1}}[h_{\varepsilon,\delta}]-\mathbb{E}_{C_{0}}[h_{\varepsilon,\delta}]}{C^{\prime}\delta^{-1}}.$
@@ 6.3 Sensitivity to Tail Dependence
Let $\delta\downarrow 0$. Since $h_{\varepsilon,\delta}\to\mathbf{1}_{B_{\varepsilon}}$ pointwise with bounded convergence, we obtain
$\lim_{\delta\downarrow 0}\mathbb{E}_{C_{i}}[h_{\varepsilon,\delta}]=\Pr_{C_{i}}\big(U\in B_{\varepsilon}\big)\qquad(i=0,1).$
Hence
$\mathrm{CSD}_{C_{2}}(C_{1})\ \geq\ c\,\delta\,\big\{\Pr_{C_{1}}(U\in B_{\varepsilon})-\Pr_{C_{0}}(U\in B_{\varepsilon})\big\}\quad\text{for all small }\delta.$
We now send $\varepsilon\downarrow 0$ while coupling $\delta=\delta(\varepsilon)$ so that $\delta(\varepsilon)\to 0$ and $\varepsilon/\delta(\varepsilon)\to 0$. This keeps the Lipschitz penalty negligible relative to the tail mass. Using the definition of $\lambda_{L}$ in A4 and the monotonicity of $t\mapsto\Pr(U\in[0,t]^{2})$,
$\frac{\Pr_{C_{i}}(U\in B_{\varepsilon})}{\varepsilon}\ \longrightarrow\ \lambda_{L}(C_{i})\qquad(i=0,1).$
Therefore for small $\varepsilon$,
$\Pr_{C_{1}}(U\in B_{\varepsilon})-\Pr_{C_{0}}(U\in B_{\varepsilon})\ =\ \big(\lambda_{L}(C_{1})-\lambda_{L}(C_{0})\big)\,\varepsilon+o(\varepsilon).$
@@
Choose the sign of the witness by replacing $h_{\varepsilon,\delta}$ with its negative if needed, so the difference above is nonnegative. Then for all sufficiently small $\varepsilon$,
$\mathrm{CSD}_{C_{2}}(C_{1})\ \geq\ c\,\delta(\varepsilon)\,\big(\lambda_{L}(C_{1})-\lambda_{L}(C_{0})\big)\,\varepsilon+o\big(\delta(\varepsilon)\,\varepsilon\big).$
Since $\lambda_{L}(C_{1})\neq\lambda_{L}(C_{0})$, the right side is strictly positive for small $\varepsilon$. Thus $\mathrm{CSD}_{C_{2}}(C_{1})>0$.
@@
Proposition 6.1 (Quantitative tail sensitivity for Clayton, $d=2$)
Let $C_{\theta}$ be the Clayton copula on $(0,1)^{2}$ with parameter $\theta>0$. For any $\theta\neq\theta^{\prime}$, there exists a constant $c(\theta,\theta^{\prime},\mathcal{H})>0$ such that
$\mathrm{CSD}_{C_{\theta^{\prime}}}(C_{\theta})\ \geq\ c(\theta,\theta^{\prime})\,|\lambda_{L}(C_{\theta})-\lambda_{L}(C_{\theta^{\prime}})|.$
Proof. Follow the proof of Theorem 5 with $C_{0}=C_{\theta^{\prime}}$ and recall that $\lambda_{L}(C_{\theta})=2^{-1/\theta}$. The Clayton score has explicit form $s(u)=\nabla\log c_{\theta^{\prime}}(u)$ with polynomial growth near the origin, so the constants in the Stein-equation construction in the proof of Theorem 3 are finite. The lower bound follows by the same localization and limiting argument with the proportionality constant depending on $\theta^{\prime}$ and the RKHS constants.
Remark 6.1 (Optional quantitative bound)
If one additionally assumes regular variation of the generator near the boundary, e.g. $\varphi_{0}(1-x)\sim ax^{\kappa}$ with $\kappa>0$, and matches the RKHS boundary rate from Theorem 3, a quantitative lower bound $\mathrm{CSD}\geq c\,|\lambda_{L}^{(1)}-\lambda_{L}^{(2)}|^{\alpha}$ can be derived. We omit this optional refinement here.
Theorem 6 (Minimax risk for the squared discrepancy)
Let $\widehat{\mathrm{CSD}}_{n}^{2}$ be any estimator of $D^{2}(Q)\equiv\mathrm{CSD}(Q,C_{0})^{2}$ based on an i.i.d. sample of size $n$ from $Q$. Work over the local nonparametric class
$\mathcal{Q}\equiv\big\{Q_{\theta}:d\mathrm{TV}(Q_{\theta},C_{0})\leq\kappa\,n^{-1/2},\ |\theta|\leq\theta_{0}\big\},$
constructed by exponential tilts along a bounded function with zero mean under $C_{0}$. Under A1A3 there exist constants $0<c\leq C<\infty$ such that
$c\,n^{-2}\ \leq\ \inf_{\widehat{\mathrm{CSD}}_{n}^{2}}\ \sup_{Q\in\mathcal{Q}}\ \mathbb{E}_{Q}\big[\big(\widehat{\mathrm{CSD}}_{n}^{2}-D^{2}(Q)\big)^{2}\big]\ \leq\ C\,n^{-2}.$
Proof. Upper bound. Your Section 6 already proves that under $Q=C_{0}$, $\mathbb{E}[\widehat{\mathrm{CSD}}_{n}^{2}]=O(n^{-1})$ and $\mathrm{Var}(\widehat{\mathrm{CSD}}_{n}^{2})=O(n^{-2})$ for the natural U statistic estimator. This gives $\mathbb{E}[(\widehat{\mathrm{CSD}}_{n}^{2}-0)^{2}]=O(n^{-2})$. The same calculation holds uniformly over a $\mathrm{TV}$ ball of radius of order $n^{-1/2}$ by a standard contiguity argument, which yields the stated $C\,n^{-2}$ bound.
@@
$\frac{dQ_{\theta}}{dC_{0}}(u)\ \equiv\ 1+\theta\,h(u).$
@@
$D(Q_{\theta})=\sup_{\|g\|_{\mathcal{H}^{d}}\leq 1}\mathbb{E}_{Q_{\theta}}[\mathcal{T}_{C_{0}}g]=\sup_{\|g\|_{\mathcal{H}^{d}}\leq 1}\mathbb{E}_{C_{0}}\big[(1+\theta h)\,\mathcal{T}_{C_{0}}g\big]=\theta\sup_{\|g\|_{\mathcal{H}^{d}}\leq 1}\mathbb{E}_{C_{0}}\big[h\,\mathcal{T}_{C_{0}}g\big],$
@@
$\inf_{\widehat{\mathrm{CSD}}_{n}^{2}}\frac{1}{2}\Big\{\mathbb{E}_{Q_{+\theta}}\big[\big(\widehat{\mathrm{CSD}}_{n}^{2}-\theta^{2}L^{2}\big)^{2}\big]+\mathbb{E}_{Q_{-\theta}}\big[\big(\widehat{\mathrm{CSD}}_{n}^{2}-\theta^{2}L^{2}\big)^{2}\big]\Big\}\ \geq\ c_{0}\,\big(D^{2}(Q_{+\theta})-D^{2}(Q_{-\theta})\big)^{2}\,\big(1-\mathrm{TV}(Q_{+\theta}^{n},Q_{-\theta}^{n})\big).$
@@
$\sup_{Q\in\{Q_{0},\,Q_{\theta}\}}\mathbb{E}_{Q}\big[\big(\widehat{\mathrm{CSD}}_{n}^{2}-D^{2}(Q)\big)^{2}\big]\ \geq\ c_{1}\,\big(D^{2}(Q_{\theta})-D^{2}(Q_{0})\big)^{2}=c_{1}\,\theta^{4}L^{4}=c_{1}\,\kappa^{4}L^{4}\,n^{-2}.$
This gives the stated $c\,n^{-2}$ lower bound with $c=c_{1}\kappa^{4}L^{4}$.
6.4 Summary of Theoretical Guarantees
  1. 1.
    Is a valid metric, as it metrizes weak convergence on the space of copula distributions (Theorem 3).
  2. 2.
    Is statistically optimal for finite samples, with its empirical estimator converging at the minimax optimal rate of $O_{P}(n^{-1/2})$ (Theorem 4, Corollary 6.1, and Theorem 6).
  3. 3.
    Is sensitive to fine-grained dependence, provably distinguishing between copulas with different tail dependence coefficients (Theorem 5).
7.  Extension Beyond Archimedean Families
Our analysis so far has leveraged the convenient generator-based form of Archimedean copulas. However, the CSD framework is far more general. The key insight is that our method does not depend on the Archimedean structure itself, but only on the analytical properties of the copula’s score function, which are captured by our Unified Regularity Conditions (Assumption 1). In this section, we formally show that all our main results apply to any copula family satisfying these conditions, thereby extending CSD’s applicability to a wide range of important models.
7.1 Generality of the CSD Framework
To extend the framework, we first note that the Copula-Stein operator can be defined for any copula $C$ with a sufficiently regular density $c$:
@@ 7.1 Generality of the CSD Framework
Theorem 7 (Zero-Mean Property for General Copulas)
Under Assumptions 1 (A1–A4), for any test function $g\in C_{c}^{1}((0,1)^{d};\mathbb{R}^{d})$:
@@
For each $j$, we apply integration by parts to the first term in the integral, integrating with respect to $u_{j}$:
@@
where $du_{-j}$ denotes integration over all coordinates except $u_{j}$. The boundary term,
$\int_{(0,1)^{d-1}}[g_{j}(u)c(u)]_{u_{j}=0}^{u_{j}=1}du_{-j}$, vanishes due to Assumption 1, which states that test functions in the RKHS vanish appropriately at the boundary.
Therefore, the integration-by-parts identity simplifies to:
@@
Substituting this back into the expression for the expectation, the two terms for each $j$ cancel out:
$\mathbb{E}_{U\sim C}[\mathcal{T}_{C_{0}}g(U)]$ $=\sum_{j=1}^{d}\left[-\int_{(0,1)^{d}}g_{j}(u)\frac{\partial c(u)}{\partial u_{j}}\,du+\int_{(0,1)^{d}}g_{j}(u)\frac{\partial c(u)}{\partial u_{j}}\,du\right]=0.$ (45)
Theorem 8 (General Copula-Stein Kernel)
Under Assumptions 1 (A1–A4), let k: $0,1]^{d}\times[0,1]^{d}\to\mathbb{R}$ be a $C^{2}$ characteristic kernel. The General Copula-Stein kernel is given by:
@@
Proof. The proof is identical to the Archimedean case (Proposition 5.2), as it only relies on the witness function representation and the reproducing property of the RKHS, which hold for any smooth copula density.
7.2 Extension of Main Theoretical Results
@@
Theorem 9 (Metrization for General Copulas)
Under Assumptions 1 (A1–A4), $k$ is characteristic and $C^{2}$, for any sequence of probability measures $(Q_{n})$ on $0,1]^{d}$:
@@
Proof. The proof follows the same structure as Theorem 3, but requires verification that the function class $\mathcal{F}_{C}=\{\mathcal{T}_{C_{0}}g:\|g\|_{\mathcal{H}^{d}}\leq 1\}$ satisfies the conditions for the Portmanteau theorem under the general regularity assumptions.
@@
Using the reproducing property (from Assumption 1), we have $|g_{j}(u)|\leq M_{k}$ and $|\partial_{u_{j}}g_{j}(u)|\leq M_{k}^{\prime}$ on any compact subset $K\Subset(0,1)^{d}$. From Assumption 1:
@@
  • Uniform continuity of $g_{j}$ and $\partial_{u_{j}}g_{j}$ (from the kernel properties in Assumption 1)
  • Uniform continuity of $s_{j}$ on compact subsets (from Assumption 1)
@@
Step 3: Convergence-Determining Property. The key observation is that the growth condition in Assumption 1 ensures that functions in $\mathcal{F}_{C}$ remain bounded on compact subsets, while the boundary condition in Assumption 1 ensures proper behavior near the boundary. The characteristic property of the kernel (from Assumption 1) then guarantees that $\mathcal{F}_{C}$ is convergence-determining.
The remainder of the proof (both directions of the equivalence) follows the same pattern as Theorem 3.
Theorem 10 (Finite-Sample Convergence for General Copulas)
Under the unified assumptions 1, let $\widehat{\text{CSD}}_{n}^{2}=\frac{1}{n^{2}}\sum_{i,j=1}^{n}k_{C}(U_{i},U_{j})$ where $U_{1},\ldots,U_{n}\sim C$ are i.i.d. Then:
@@
Proof. Step 1: Mean Square Error Analysis. The decomposition into diagonal and off-diagonal terms proceeds exactly as in Theorem 4:
@@
For $i\neq j$, by independence and the zero-mean property (Theorem 7):
@@
Step 3: Concentration Inequality. The concentration bound follows from the same U-statistic theory, with the uniform boundedness now guaranteed by Assumptions C1-C2 on compact subsets where most of the probability mass lies.
7.3 Examples of Non-Archimedean Copulas

### v2

@@ Appendix A Proofs
Since $\mathbb{E}[\widehat{\mathrm{CSD}}_{n}^{2}]=O(n^{-1})$, it follows that $\widehat{\mathrm{CSD}}_{n}^{2}=O_{P}(n^{-1})$. Finally, taking square roots (a continuous mapping) yields $\widehat{\mathrm{CSD}}_{n}=O_{P}(n^{-1/2})$. ∎
A.10 Proof of Theorem 6.5
Proof.
Write $C_{0}\equiv C_{2}$. We prove the lower-tail case; the upper tail follows by symmetry. Fix $\varepsilon\in(0,1)$ and let $B_{\varepsilon}=[0,\varepsilon]^{2}$. Let $\rho_{\delta}$ be a standard $C^{\infty}$ mollifier supported in a ball of radius $\delta$, and define
[…]
By Propositions 5.1–5.2, $\mathrm{CSD}_{C_{0}}(C_{1})=\mathrm{MMD}_{k_{C_{0}}}(C_{1},C_{0})$, hence
$\mathrm{CSD}_{C_{0}}(C_{1})=\sup_{\|f\|_{\mathcal{H}_{k_{C_{0}}}}\leq 1}\Big(\mathbb{E}_{C_{1}}[f(U)]-\mathbb{E}_{C_{0}}[f(U)]\Big).$ (73)
Under A3, $k_{C_{0}}$ is $c$-universal on $0,1]^{2}$, so $\mathcal{H}_{k_{C_{0}}}$ is dense in $C([0,1]^{2})$. Since $\bar{h}_{\varepsilon,\delta}$ is continuous and bounded, for any $\eta>0$ there exists $f_{\varepsilon,\delta}\in\mathcal{H}_{k_{C_{0}}}$ such that $\|f_{\varepsilon,\delta}-\bar{h}_{\varepsilon,\delta}\|_{\infty}\leq\eta$. Therefore,
$\mathrm{CSD}_{C_{0}}(C_{1})\;\geq\;\frac{\mathbb{E}_{C_{1}}[f_{\varepsilon,\delta}]-\mathbb{E}_{C_{0}}[f_{\varepsilon,\delta}]}{\|f_{\varepsilon,\delta}\|_{\mathcal{H}_{k_{C_{0}}}}}\;\geq\;\frac{\mathbb{E}_{C_{1}}[\bar{h}_{\varepsilon,\delta}]-\mathbb{E}_{C_{0}}[\bar{h}_{\varepsilon,\delta}]-2\eta}{\|f_{\varepsilon,\delta}\|_{\mathcal{H}_{k_{C_{0}}}}}.$ (74)
@@
Since $\mathbb{E}_{C_{0}}[\bar{h}_{\varepsilon,\delta}]=0$, we have $\mathbb{E}_{C_{1}}[\bar{h}_{\varepsilon,\delta}]=\mathbb{E}_{C_{1}}[h_{\varepsilon,\delta}]-\mathbb{E}_{C_{0}}[h_{\varepsilon,\delta}]$. Letting $\delta\downarrow 0$ and using bounded convergence yields
$\lim_{\delta\downarrow 0}\Big(\mathbb{E}_{C_{i}}[h_{\varepsilon,\delta}]\Big)=\Pr_{C_{i}}(U\in B_{\varepsilon}),\qquad i\in\{0,1\}.$ (75)
Thus for $\delta$ small,
$\mathbb{E}_{C_{1}}[h_{\varepsilon,\delta}]-\mathbb{E}_{C_{0}}[h_{\varepsilon,\delta}]=\Pr_{C_{1}}(U\in B_{\varepsilon})-\Pr_{C_{0}}(U\in B_{\varepsilon})+o(1).$ (76)
@@
By the definition of $\lambda_{L}$ (Assumption A4), $\Pr_{C_{i}}(U\in B_{\varepsilon})=\lambda_{L}(C_{i})\varepsilon+o(\varepsilon)$ as $\varepsilon\downarrow 0$, so if $\lambda_{L}(C_{1})\neq\lambda_{L}(C_{0})$, then for all sufficiently small $\varepsilon$ the difference $\Pr_{C_{1}}(U\in B_{\varepsilon})-\Pr_{C_{0}}(U\in B_{\varepsilon})$ has a fixed nonzero sign. Choose the sign by replacing $h_{\varepsilon,\delta}$ with $h_{\varepsilon,\delta}$ if needed. Then pick $\eta>0$ small enough so the numerator above is strictly positive. Hence $\mathrm{CSD}_{C_{0}}(C_{1})>0$.
@@
A.11 Proof of Theorem 6.6
Proof.
Upper bound. Your Section 6 already proves that under $Q=C_{0}$, $\mathbb{E}[\widehat{\mathrm{CSD}}_{n}^{2}]=O(n^{-1})$ and $\mathrm{Var}(\widehat{\mathrm{CSD}}_{n}^{2})=O(n^{-2})$ for the natural U statistic estimator. This gives $\mathbb{E}[(\widehat{\mathrm{CSD}}_{n}^{2}-0)^{2}]=O(n^{-2})$. The same calculation holds uniformly over a $\mathrm{TV}$ ball of radius of order $n^{-1/2}$ by a standard contiguity argument, which yields the stated $C\,n^{-2}$ bound.
@@
$\frac{dQ_{\theta}}{dC_{0}}(u)\ \equiv\ 1+\theta\,h(u).$ (77)
Note that $Q_{\theta}$ need not preserve uniform margins; this lower bound is stated for a TV neighborhood in $\mathcal{P}((0,1)^{d})$ rather than a copula-restricted subclass.
@@
$D(Q_{\theta})=\sup_{\|g\|_{\mathcal{H}^{d}}\leq 1}\mathbb{E}_{Q_{\theta}}[\mathcal{T}_{c_{0}}g]=\sup_{\|g\|_{\mathcal{H}^{d}}\leq 1}\mathbb{E}_{C_{0}}\big[(1+\theta h)\,\mathcal{T}_{c_{0}}g\big]=\theta\sup_{\|g\|_{\mathcal{H}^{d}}\leq 1}\mathbb{E}_{C_{0}}\big[h\,\mathcal{T}_{c_{0}}g\big],$ (78)
@@
$\inf_{\widehat{\mathrm{CSD}}_{n}^{2}}\frac{1}{2}\Big\{\mathbb{E}_{Q_{+\theta}}\big[\big(\widehat{\mathrm{CSD}}_{n}^{2}-\theta^{2}L^{2}\big)^{2}\big]+\mathbb{E}_{Q_{-\theta}}\big[\big(\widehat{\mathrm{CSD}}_{n}^{2}-\theta^{2}L^{2}\big)^{2}\big]\Big\}\ \geq\ c_{0}\,\big(D^{2}(Q_{+\theta})-D^{2}(Q_{-\theta})\big)^{2}\,\big(1-\mathrm{TV}(Q_{+\theta}^{n},Q_{-\theta}^{n})\big).$ (80)
@@
$\sup_{Q\in\{Q_{0},\,Q_{\theta}\}}\mathbb{E}_{Q}\big[\big(\widehat{\mathrm{CSD}}_{n}^{2}-D^{2}(Q)\big)^{2}\big]\ \geq\ c_{1}\,\big(D^{2}(Q_{\theta})-D^{2}(Q_{0})\big)^{2}=c_{1}\,\theta^{4}L^{4}=c_{1}\,\kappa^{4}L^{4}\,n^{-2}.$ (81)
This gives the stated $c\,n^{-2}$ lower bound with $c=c_{1}\kappa^{4}L^{4}$. ∎
Appendix B Extension Beyond Archimedean Families
Our analysis so far has leveraged the convenient generator-based form of Archimedean copulas. However, the CSD framework is far more general. The key insight is that our method does not depend on the Archimedean structure itself, but only on the analytical properties of the copula’s score function, which are captured by our Unified Regularity Conditions (Assumption 4.2). In this section, we formally show that all our main results apply to any copula family satisfying these conditions, thereby extending CSD’s applicability to a wide range of important models.
B.1 Generality of the CSD Framework
To extend the framework, we first note that the Copula–Stein operator can be defined for any copula $C$ with a sufficiently regular density $c$:
@@ B.1 Generality of the CSD Framework
Theorem B.1 (Zero-Mean Property for General Copulas).
Under Assumptions 4.2 (A1–A4), for any test function $g\in C_{c}^{1}((0,1)^{d};\mathbb{R}^{d})$,
@@
Fix $j$. Applying integration by parts with respect to $u_{j}$ yields
@@
where $du_{-j}$ denotes integration over all coordinates except $u_{j}$. Since $g\in C_{c}^{1}((0,1)^{d};\mathbb{R}^{d})$ has compact support contained in $(0,1)^{d}$, we have $g_{j}(u)=0$ in a neighborhood of the boundary, and thus the boundary term vanishes. Therefore,
@@
Substituting back, the two terms cancel for each $j$, giving $\mathbb{E}_{U\sim C}[\mathcal{T}_{c_{0}}g(U)]=0$. ∎
Theorem B.2 (General Copula-Stein Kernel).
Under Assumptions 4.2 (A1–A4), let k: $0,1]^{d}\times[0,1]^{d}\to\mathbb{R}$ be a $C^{2}$ characteristic kernel. The General Copula-Stein kernel is given by:
@@
Proof.
The proof is identical to the Archimedean case (Proposition 5.6), as it only relies on the witness function representation and the reproducing property of the RKHS, which hold for any smooth copula density. ∎
B.2 Extension of Main Theoretical Results
@@
Theorem B.3 (Metrization for General Copulas).
Under Assumptions 4.2 (A1–A4), $k$ is characteristic and $C^{2}$, for any sequence of probability measures $(Q_{n})$ on $0,1]^{d}$:
@@
Proof.
The proof follows the same structure as Theorem 6.2, but requires verification that the function class $\mathcal{F}_{C}=\{\mathcal{T}_{c_{0}}g:\|g\|_{\mathcal{H}^{d}}\leq 1\}$ satisfies the conditions for the Portmanteau theorem under the general regularity assumptions.
@@
Using the reproducing property (from Assumption 4.2), we have $|g_{j}(u)|\leq M_{k}$ and $|\partial_{u_{j}}g_{j}(u)|\leq M_{k}^{\prime}$ on any compact subset $K\Subset(0,1)^{d}$. From Assumption 4.2:
@@
  • Uniform continuity of $g_{j}$ and $\partial_{u_{j}}g_{j}$ (from the kernel properties in Assumption 4.2)
  • Uniform continuity of $s_{j}$ on compact subsets (from Assumption 4.2)
@@
Step 3: Convergence-Determining Property. The key observation is that the growth condition in Assumption 4.2 ensures that functions in $\mathcal{F}_{C}$ remain bounded on compact subsets, while the boundary condition in Assumption 4.2 ensures proper behavior near the boundary. The characteristic property of the kernel (from Assumption 4.2) then guarantees that $\mathcal{F}_{C}$ is convergence-determining.
The remainder of the proof (both directions of the equivalence) follows the same pattern as Theorem 6.2. ∎
Theorem B.4 (Finite-Sample Convergence for General Copulas).
Under the unified assumptions 4.2, let $\widehat{\text{CSD}}_{n}^{2}=\frac{1}{n^{2}}\sum_{i,j=1}^{n}k_{C}(U_{i},U_{j})$ where $U_{1},\ldots,U_{n}\sim C$ are i.i.d. Then:
@@
Proof.
Step 1: Mean Square Error Analysis. The decomposition into diagonal and off-diagonal terms proceeds exactly as in Theorem 6.3:
@@
For $i\neq j$, by independence and the zero-mean property (Theorem B.1):
@@
Step 3: Concentration Inequality. The concentration bound follows from the same U-statistic theory, with the uniform boundedness now guaranteed by Assumptions C1-C2 on compact subsets where most of the probability mass lies. ∎
B.3 Examples of Non-Archimedean Copulas